50.0 mL of 0.090 M nitrous acid (HNO2, Ka = 7.1 x 10-4), is titrated with 0.100 M NaOH, requiring 45.0 mL of strong base to reach the equivalence point.

(a) What will be the pH after 35.0 mL of NaOH have been added?

(b) What will be the pH at the equivalence point?

(c) What will be the pH after 60.0 mL of NaOH have been added?
Answer #1

Titration of Weach a aud with • o neutralisation reaction takes place between nitrous and (HMO2) and strong base odium hy dro- pH = 3.1487 + 0.544 - pH = 3.6927 - pH of solution after addition of 3 of ma oH = 3.6927 Nach in 45.0m2 of o.100 m Solution